### v3

@@ 4 Mutual locality and Mirror model
The solutions to equation (33) can be investigated by analyzing the construction of the spectral flow of the mirror orbifold Parkhomenko:2024mxq ; Parkhomenko:2022kju . However, the remarkable fact is that the field space of the mirror orbifold arose automatically as soon as we constructed the state space of the original orbifold.
To verify this, let us consider the mirror version of (23). It uses so-called mirror spectral flow twisting, which starts with the anti-chiral primary field so the mirror realization of primary states is given by
@@
$\tilde{\Psi}^{\vec{l},\vec{\bar{l}}}_{\vec{t}+\vec{w}^{*},\vec{\bar{t}}}(z,\bar{z})=\tilde{V}_{\vec{t}+\vec{w}^{*}}^{\vec{l}}(z)\bar{V}_{\vec{\bar{t}}}^{\vec{l}}(\bar{z}),$ (39)
$\tilde{V}_{\vec{t}+\vec{w}^{*}}^{\vec{l}}(z)=\prod_{i=1}^{r}\tilde{V}_{t_{i}+w_{i}^{*}}^{l_{i}}(z),$
$\tilde{V}_{t_{i}+w_{i}^{*}}^{l_{i}}(z)=\left\{\begin{array}[]{l}\left(UG_{-\frac{1}{2}}^{-}\right)_{i}^{t_{i}+w_{i}^{*}}\Phi^{l_{i}}_{c}(z),\quad w_{i}^{*}=l_{i}-2t_{i}-w_{i},\\ \left(UG_{-\frac{1}{2}}^{-}\right)_{i}^{t_{i}+w_{i}^{*}-l_{i}-1}U_{i}\left(UG_{-\frac{1}{2}}^{-}\right)_{i}^{l_{i}}\Phi^{l_{i}}_{c}(z),\quad w_{i}^{*}=k_{i}+2+l_{i}-2t_{i}-w_{i}\simeq l_{i}-2t_{i}-w_{i}.\end{array}\right$.
The only difference is that this field appears as twisted by the element of another group
$\vec{w}^{*}=\vec{q}-\vec{w}\in G_{\text{adm}}^{*}.$ (40)
The vectors (40) form a group indeed, since with respect to the operator algebra of the orbifold model the vectors $\vec{q}$ are conserved while the vectors $\vec{w}$ are additive. It is clear also that the group $G_{\text{adm}}^{*}$ is admissible since the primary fields, which correspond to $(3,0)$ and $(0,3)$ differential forms on CY manifold are among the set of mutually local fields. Hence, (39) can be interpreted as a field in the orbifold model that is built by the dual group:
$\tilde{\Psi}^{\vec{l},\vec{\bar{l}}}_{\vec{t}+\vec{w}^{*},\vec{\bar{t}}}(z,\bar{z})\in M_{\vec{k}}/G_{\text{adm}}^{*}.$ (41)
This orbifold model is isomorphic to $M_{\vec{k}}/G_{\text{adm}}$ by the construction. Moreover, as we will see shortly this model is mirror to the initial orbifold with group $G_{\text{adm}}$.
@@
$\sum_{i=1}^{r}\frac{w_{1i}w_{2i}^{*}+w_{2i}w_{1i}^{*}}{k_{i}+2}\in\mathbb{Z}.$ (43)
@@ 4.1 Dual group
In the case where the composite model consists of minimal models of arbitrary types, equations (44) can be considered as defining the generalization of the Berglund-Hubsch-Krawitz dual group. More precisely, we define the dual group as
@@
We have thereby demonstrated that both the definition of the dual group and the construction of the dual orbifold model (mirror model) naturally extend to composite models with an arbitrary modular invariant.
@@ 4.2 Mirror Symmetry of mutually dual orbifolds
Its holomorphic factor may be off-diagonal in $l_{i}$, but has to be chiral primary state also. It is not difficult to see that this happens if
@@
Besides, these relations have been obtained under the assumption $q_{i}\leq l_{i}$. If the charge exceeds $q_{i}>l_{i}$ we must identify the fields as in (8). The charge becomes $\tilde{q}_{i}=q_{i}+(k_{i}+2)$, but this does not affect the condition on the dual twist, since $w_{i}^{*}=q_{i}-w_{i}\simeq q_{i}+(k_{i}+2)-w_{i}=\tilde{q}_{i}-w_{i}$. Then the conditions (50),(51) are rewritten with a replacement $\tilde{l}_{i}=k-l_{i}$, and $\tilde{q}_{i}=\tilde{l}_{i}-2\tilde{t}_{i}$.
@@ 5 Examples of orbifold models with a non-diagonal modular invariants
where the vector $(2,2,2,2)$ is the generator of the first group, while the vectors $(2,2,2,2)$, $(1,0,0,12)$, $(1,2,4,6)$, $(0,0,4,14)$ are the generators of the second group. For this pair of models we write out the $(a,c)$-fields with the formulas (23)-(26) in mind.
@@
where $x_{i}\in\mathbb{P}^{3}$, and $y_{i}\in\mathbb{P}^{2}$. The Hödge numbers are $h^{2,1}(X)=h^{1,2}(X)=35,\quad h^{2,2}(X)=h^{1,1}(X)=8$, $h^{0,0}(X)=h^{0,3}(X)=h^{3,0}(X)=h^{3,3}(X)=1$. This orbifold model has been considered by D.Gepner Gepner:1987hi as an initial step in his construction of three-generation model of superstring compactification.
The $(a,c)$-fields of the mirror orbifold obviously coincide with De Rham cohomology of the mirror Calabi-Yau manifold, which is nothing but the orbifold of $X$ by ”maximal admissible” group $G^{*}$.
@@
The structure of the tables is as follows. The first two columns of the tables contain the number of $(a,c)$ fields with charge $(Q_{L},Q_{R})$ and corresponding De-Rham cohomologies of CY. The untwisted fields from the composite model, which are located in the penultimate column, have the form $\Phi^{\vec{l}}_{\vec{q}}\bar{\Phi}^{\vec{\bar{l}}}_{c}$. The set of charges $\vec{q}$ of a holomorphic state coincides with the set of charges of an anti-holomorphic chiral state, that is $q_{i}=\bar{l}_{i}$. In the case when $q_{i}>l_{i}$ the corresponding field should be identified according to the (8). In the ”Field” column are located the $(a,c)$ fields in the orbifold model, twisting by the element $\vec{w}$ from the last column.
@@
We write out again the $(a,c)$-fields of the first orbifold and relate them with De Rham cohomology groups of the Calabi-Yau manifold $Y$, which is defined as an orbifold of the intersection of hypersurfaces (53) by the group $\tilde{G}$ in $\mathbb{P}^{2}\times\mathbb{P}^{3}/\tilde{G}$: The Hödge numbers are $h^{2,1}(Y)=h^{1,2}(Y)=23,\quad h^{2,2}(Y)=h^{1,1}(Y)=14$, $h^{0,0}(Y)=h^{0,3}(Y)=h^{3,0}(Y)=h^{3,3}(Y)=1$.
The $(a,c)$-fields of the mirror orbifold correspond to De Rham cohomology of the mirror Calabi-Yau manifold which is defined as the orbifold of the original one (53) by dual group $\tilde{G}^{*}$.
@@
As we see, every orbifold model contains holomorphic and anti-holomorphic currents of charge 3, corresponding to the nowhere-vanishing (3,0) and (0,3) forms on the Calabi-Yau manifold. The mutual locality of these currents with the other fields of the model enforces the admissible group constraint (17) introduced in Section 3.
@@ 6 Conclusion
We presented the explicit construction of fields in orbifolds of products of $N=(2,2)$ minimal models of ADE types, using spectral flow twisting and conformal bootstrap axioms. The spaces of fields of the initial orbifold model $M_{\vec{k}}/G_{\text{adm}}$ and their mirror $M_{\vec{k}}/G_{\text{adm}}^{*}$ are shown to be isomorphic, with the groups $G_{\text{adm}}$ and $G_{\text{adm}}^{*}$ related by Berglund-Hubsch-Krawitz-like condition. We shown that the spaces of fields of the initial orbifold $M_{\vec{k}}/G_{\text{adm}}$ and its mirror $M_{\vec{k}}/G_{\text{adm}}^{*}$ appear in the construction simultaneously. In other words, the mirror symmetry isomorphism of fields is built into our conformal bootstrap construction, as was suggested initially in Lerche:1989uy . Our results can also be considered as a conformal bootstrap version of the modular bootstrap construction of mirror orbifolds from GrPl . The geometric interpretation of the orbifolds of ADE models that we consider was studied in Fuchs:1989pt ; Lynker , where it was shown that such orbifolds are related to Calabi-Yau manifolds defined as a complete intersection in projective spaces. In this paper, we thereby have given an explicit construction of fields in $\sigma$-models on the Calabi-Yau manifolds studied there.
@@ Acknowledgments
We are grateful to prof. A. Belavin, and to A. Litvinov for useful discussions. The work of B. Eremin was performed in Kharkevich IITP within the framework of the state assignment. The work of S. Parkhomenko has been supported by the Russian Science Foundation grant 23-12-00333.
@@ Appendix A Tables
Term in $Z_{E_{7}}$ Field: $\Phi_{q}^{l}\bar{\Phi}_{\bar{q}}^{\bar{l}}$ Area of $q$

### v4

@@ 4 Mutual locality and Mirror model
The solutions to equation (40) can be investigated by analyzing the construction of the spectral flow of the dual orbifold Parkhomenko:2024mxq ; Parkhomenko:2022kju (whose notion will become clear soon). However, the remarkable fact is that the field space of the mirror orbifold arose automatically as soon as we constructed the state space of the original orbifold.
To verify this, let us consider the mirror version of (30), (40) Parkhomenko:2022kju . It uses so-called mirror spectral flow twisting, which starts with the anti-chiral primary field so the mirror realization of a Minimal model primary states is given by
@@
$\Psi^{\vec{l},\vec{\bar{l}}}_{\vec{t}+\vec{w}^{*},\vec{\bar{t}}}(z,\bar{z})=V_{\vec{t}+\vec{w}^{*}}^{\vec{l}}(z)\bar{V}_{\vec{\bar{t}}}^{\vec{l}}(\bar{z}),$ (46)
$V_{\vec{t}+\vec{w}^{*}}^{\vec{l}}(z)=\prod_{i=1}^{r}V_{t_{i}+w_{i}^{*}}^{l_{i}}(z),$
$V_{t_{i}+w_{i}^{*}}^{l_{i}}(z)=\left\{\begin{array}[]{l}\left(UG_{-\frac{1}{2}}^{-}\right)_{i}^{t_{i}+w_{i}^{*}}\Phi^{l_{i}}_{c}(z),\quad w_{i}^{*}=l_{i}-2t_{i}-w_{i},\\ \left(UG_{-\frac{1}{2}}^{-}\right)_{i}^{t_{i}+w_{i}^{*}-l_{i}-1}U_{i}\left(UG_{-\frac{1}{2}}^{-}\right)_{i}^{l_{i}}\Phi^{l_{i}}_{c}(z),\quad w_{i}^{*}=k_{i}+2+l_{i}-2t_{i}-w_{i}\simeq l_{i}-2t_{i}-w_{i}.\end{array}\right$.
The only difference is that this field appears as twisted by the element of another group, denoted $\check{G}$. This group is defined as follows:
$\check{G}=\left\{\vec{w}^{*}=\vec{l}-2\vec{t}-\vec{w}=\vec{q}-\vec{w}\right\}.$ (47)
Recall that here, the vectors $\vec{q}$ label all possible charges of fields in composite model $M_{\vec{k}}$ that lie in the twisted sector of the element $\vec{w}$ of the orbifold $M_{\vec{k}}/G_{\text{adm}}$. Thus, these fields satisfy the mutual locality condition (40).
The set of vectors in (47) indeed forms a group, as the vectors $\vec{w}^{*}$ are additive under the operator algebra of the orbifold model. This additivity stems from the fact that the $\vec{q}$ correspond to conserved $U(1)$ charges of the initial composite model $M_{\vec{k}}$ (i.e charges of two fields add under fusion), while the vectors $\vec{w}$ are additive by definition. It is also clear, that the elements $\vec{w}^{*}$ obey (22) and hence, the group $\check{G}$ is admissible because the primary fields, which correspond to $(3,0)$ and $(0,3)$ differential forms on CY manifold are among the set of mutually local fields. Since this is an admissible group, we will refer to it as $\check{G}\equiv G_{\text{adm}}^{*}$. Hence, (46) can be interpreted as a field in the orbifold model that is built by the dual group:
$\Psi^{\vec{l},\vec{\bar{l}}}_{\vec{t}+\vec{w}^{*},\vec{\bar{t}}}(z,\bar{z})\in M_{\vec{k}}/G_{\text{adm}}^{*}.$ (48)
This dual orbifold model is isomorphic to $M_{\vec{k}}/G_{\text{adm}}$ by the construction. Moreover, as we will see shortly, this model is mirror to the initial orbifold with group $G_{\text{adm}}$.
@@
$\sum_{i=1}^{r}\frac{w_{1i}w_{2i}^{*}+w_{2i}w_{1i}^{*}}{k_{i}+2}\in\mathbb{Z},\quad\vec{w}_{1,2}\in G_{adm},\quad\vec{w}^{*}_{1,2}\in G^{*}_{adm}.$ (50)
@@ 4.1 Dual group
In the case where the composite model consists of minimal models of arbitrary types, equations (51) can be considered as defining the generalization of the Berglund-Hubsch-Krawitz dual group. More precisely, according to our discussion above, we define the dual group as
@@
In other words, the set of mutually local fields of the model $M_{\vec{k}}/G_{\text{adm}}$, which appear in the twisted sector $\vec{w}$ is given by the elements $\vec{w}^{*}\in G_{\text{adm}}^{*}$. From the other hand, the set of mutually local fields of the model $M_{\vec{k}}/G_{\text{adm}}^{*}$, which appear in the twisted sector $\vec{w}^{*}$ is given by the elements $\vec{w}\in G_{\text{adm}}$.
@@
We have thereby demonstrated that both the definition of the dual group and the construction of the dual orbifold model (mirror model) naturally extend to composite models with an arbitrary ADE modular invariant.
@@ 4.2 Mirror Symmetry of mutually dual orbifolds
Its holomorphic factor may be off-diagonal in $l_{i}$, but has to be chiral primary state also. It is not difficult to see from (30) - (31), that this happens if
@@
Besides, these relations have been obtained under the assumption $|q_{i}|\leq l_{i}$. If the charge exceeds $|q_{i}|>l_{i}$ we must identify the fields as in (12). The charge becomes $\tilde{q}_{i}=q_{i}+(k_{i}+2)$, but this does not affect the condition on the dual twist, since $w_{i}^{*}=q_{i}-w_{i}\simeq q_{i}+(k_{i}+2)-w_{i}=\tilde{q}_{i}-w_{i}$. Then the conditions (57),(58) are rewritten with a replacement $\tilde{l}_{i}=k-l_{i}$, and $\tilde{q}_{i}=\tilde{l}_{i}-2\tilde{t}_{i}$.
@@ 5 Examples of orbifold models with a non-diagonal modular invariants
where the vector $(2,2,2,2)$ is the generator of the first group, while the vectors $(2,2,2,2)$, $(1,0,0,12)$, $(1,2,4,6)$, $(0,0,4,14)$ are the generators of the second group. For this pair of models we write out the $(a,c)$-fields which are realized by the formulas (30)-(33).
@@
where $x_{i}\in\mathbb{P}^{3}$, and $y_{i}\in\mathbb{P}^{2}$. The Hodge numbers are $h^{2,1}(X)=h^{1,2}(X)=35,\quad h^{2,2}(X)=h^{1,1}(X)=8$, $h^{0,0}(X)=h^{0,3}(X)=h^{3,0}(X)=h^{3,3}(X)=1$. This orbifold model has been considered by D.Gepner Gepner:1987hi as an initial step in his construction of three-generation model of superstring compactification.
The $(a,c)$-fields of the mirror orbifold obviously coincide with De Rham cohomology of the mirror Calabi-Yau manifold, which is nothing but the orbifold of $X$ by ”maximal admissible” group $G^{*}$ and at the same time, they are $(c,c)$-fields in the initial orbifold model, as can be seen directly from our construction.
@@
The structure of the tables is as follows. The first two columns of the tables contain the number of $(a,c)$ fields with charge $(Q_{L},Q_{R})$ and corresponding De-Rham cohomologies of CY. The untwisted fields from the composite model, which are located in the penultimate column, have the form $\Phi^{\vec{l}}_{\vec{q}}\bar{\Phi}^{\vec{\bar{l}}}_{c}$. The set of charges $\vec{q}$ of a holomorphic state coincides with the set of charges of an anti-holomorphic chiral state, that is $q_{i}=\bar{l}_{i}$. In the case when $|q_{i}|>l_{i}$ the corresponding field should be identified according to the (12). In the ”Field” column are located the $(a,c)$ fields in the orbifold model, twisting by the element $\vec{w}$ from the last column.
@@
We write out again the $(a,c)$-fields of the first orbifold and relate them with De Rham cohomology groups of the Calabi-Yau manifold $Y$, which is defined as an orbifold of the intersection of hypersurfaces (60) by the group $\tilde{G}$ in $\mathbb{P}^{2}\times\mathbb{P}^{3}/\tilde{G}$: The Hodge numbers are $h^{2,1}(Y)=h^{1,2}(Y)=23,\quad h^{2,2}(Y)=h^{1,1}(Y)=14$, $h^{0,0}(Y)=h^{0,3}(Y)=h^{3,0}(Y)=h^{3,3}(Y)=1$.
The $(a,c)$-fields of the mirror orbifold correspond to De Rham cohomology of the mirror Calabi-Yau manifold which is defined as the orbifold of the original one (60) by dual group $\tilde{G}^{*}$. From the other hand, as can be seen directly from our construction, these fields consist of the $(c,c)$ fields of $\tilde{G}$ orbifold.
@@
As we can see, every orbifold model we have considered here, contains holomorphic and anti-holomorphic currents of charge 3, corresponding to the nowhere-vanishing (3,0) and (0,3) forms on the Calabi-Yau manifold. As can be directly checked, the mutual locality of these currents with the other fields of the model enforces the admissible group constraint (23) introduced in Section 3. It can be shown that this is a general result: for orbifold model $M_{\vec{k}}/G_{\text{adm}}$ composed of minimal models $M_{k_{i}}$ with modular invariants from ADE list, the mutual locality of the fields with charges $(\pm 3,0)$ and $(0,\pm 3)$ with all other fields in the orbifold model is equivalent to condition (23).
@@ 6 Conclusion
We presented the explicit construction of fields in orbifolds of products of $N=(2,2)$ minimal models of ADE types, using spectral flow twisting and conformal bootstrap axioms. The spaces of fields of the initial orbifold model $M_{\vec{k}}/G_{\text{adm}}$ and their mirror $M_{\vec{k}}/G_{\text{adm}}^{*}$ are shown to be isomorphic, with the groups $G_{\text{adm}}$ and $G_{\text{adm}}^{*}$ related by Berglund-Hubsch-Krawitz-like condition. We shown that the spaces of fields of the initial orbifold $M_{\vec{k}}/G_{\text{adm}}$ and its mirror $M_{\vec{k}}/G_{\text{adm}}^{*}$ appear in the construction simultaneously. In other words, the mirror symmetry isomorphism of fields is built into the conformal bootstrap construction, as was suggested initially in Lerche:1989uy . Our results can also be considered as a conformal bootstrap version of the modular bootstrap construction of mirror orbifolds from GrPl . The geometric interpretation of the orbifolds of ADE models that we consider was studied in Fuchs:1989pt ; Lynker , where it was shown that such orbifolds are related to Calabi-Yau manifolds defined as a complete intersection in projective spaces. In this paper, we thereby have given an explicit construction of fields in $\sigma$-models on the Calabi-Yau manifolds studied there.
@@
A natural extension of this work would be to generalize our explicit construction of fields to the orbifolds of products of minimal models with the broader class of modular invariants found in Gannon:1996hp .
@@ Acknowledgments
We are grateful to prof. A. Belavin, and to A. Litvinov for useful discussions. The work of B. Eremin was carried out within the state assignment of Ministry of Science and Higher Education of the Russian Federation for IITP RAS. The work of S. Parkhomenko was carried out at Landau Institute for Theoretical Physics in the framework of the state assignment FFWR-2024-0012.
@@ Appendix A Tables
Term in $Z_{E_{7}}$ Field: $\Phi_{q}^{l}\bar{\Phi}_{\bar{q}}^{\bar{l}}$ Range of $q$ (only even values)